# Disentangled Representation Learning for Parametric Partial Differential Equations

### v1

@@ Abstract
Neural operators (NOs) have demonstrated remarkable success in learning mappings between function spaces, serving as efficient approximators for the forward solutions of complex physical systems governed by partial differential equations (PDEs). However, while effective as black-box solvers, they offer limited insight into the underlying physical mechanism, due to the lack of interpretable representations of the physical parameters that drive the system. To tackle this challenge, we propose a new paradigm for learning disentangled representations from neural operator parameters, thereby effectively solving an inverse problem. Specifically, we introduce DisentangO, a novel hyper-neural operator architecture designed to unveil and disentangle the latent physical factors of variation embedded within the black-box neural operator parameters. At the core of DisentangO is a multi-task neural operator architecture that distills the varying parameters of the governing PDE through a task-wise adaptive layer, coupled with a hierarchical variational autoencoder that disentangles these variations into identifiable latent factors. By learning these disentangled representations, our model not only enhances physical interpretability but also enables more robust generalization across diverse physical systems. Empirical evaluations across supervised, semi-supervised, and unsupervised learning contexts show that DisentangO effectively extracts meaningful and interpretable latent features, bridging the divide between predictive performance and physical understanding in neural operator frameworks.
@@ 1 Introduction
Interpretability in machine learning (ML) refers to the ability to understand and explain how models make decisions (Rudin et al., 2022; Molnar, 2020). As ML systems grow more complex, especially with the use of deep learning and ensemble methods (Sagi and Rokach, 2018), the reasoning behind their predictions can become opaque. Interpretability addresses this challenge by making the models more transparent, enabling users to trust the outcomes, detect biases, and identify potential flaws. It is a critical factor in applying AI responsibly, especially in fields where accountability and fairness are essential (Cooper et al., 2022). In physics, where models endeavor to capture the governing laws and physical principles, understanding how a model arrives at its predictions is vital for verifying that it aligns with known scientific theories. Interpretable models also allow researchers to gain insights into the learned relationships, ensuring that the predictions are physically meaningful and consistent with real-world behaviors. This transparency is key for advancing scientific discovery, validating results, and enhancing trust in models used for complex physical systems.
@@
Discovering interpretable representations of physical parameters in learning physical systems is challenging due to the intricate nature of real-world phenomena and the often implicit relationships between variables. In physics, parameters like force, energy, and velocity are governed by well-established laws, and extracting these parameters in a way that aligns with physical intuition requires models that go beyond mere pattern recognition. Traditional ML models may fit the data but fail to represent these parameters in a physically interpretable way. To address this, recent developments include the integration of physical constraints into the learning process (Raissi et al., 2019), such as embedding conservation laws (Liu et al., 2023, 2024a), symmetries (Mattheakis et al., 2019), or invariances (Cohen and Welling, 2016) directly into model architectures. Additionally, methods like symbolic regression (Biggio et al., 2021) and sparse modeling (Carroll et al., 2009) aim to discover simple, interpretable expressions that capture the underlying dynamics. However, balancing model complexity with interpretability remains a significant challenge, as overly complex models may obscure the true physical relationships, while oversimplified models risk losing critical details of the system’s behavior.
@@
In this work, we introduce DisentangO, a novel variational hierarchical neural operator architecture to disentangle physical factors of variation from black-box neural operator parameters for solving parametric PDEs. Neural operators (NOs) (Li et al., 2020a, c) learn mappings between infinite-dimensional function spaces in the form of integral operators, making them powerful tools for discovering continuum physical laws by manifesting the mappings between spatial and/or spatiotemporal data; see You et al. (2022a); Liu et al. (2024a, b, 2023); Ong et al. (2022); Cao (2021); Lu et al. (2019, 2021); Goswami et al. (2022); Gupta et al. (2021) and references therein. However, most NOs play the role of efficient forward surrogates for the underlying physical system under a supervised learning setting. As a result, they act as black-box universal approximators for a single physical system, and are lack of interpretability with respect to the underlying physical laws. In contrast, the key innovation of DisentangO lies in its use of a hierarchical architecture that distills the varying physical parameters of the governing PDE from multiple physical systems through an unsupervised learning setting. The distilled variations are further disentangled into distinct physical factors to enhance physical understanding and promote robust generalization. Consequently, DisentangO effectively extracts meaningful and interpretable physical features, thereby simultaneously solving both the forward and inverse problems. Our key contributions are:
@@
  • We propose a novel variational hierarchical neural operator architecture, which we coin DisentangO. DisentangO extracts the key physical factors of variation from black-box neural operator parameters of multiple physical systems. These factors are then disentangled into distinct latent factors that enhance physical interpretation and robust generalization.
@@
  • We explore the practical utility of disentanglement rather than pursuing it as an end goal, and perform experiments across a broad range of settings including supervised, semi-supervised, and unsupervised learning. Our results show that DisentangO effectively extracts meaningful and interpretable physical features for enhanced interpretability.
@@ 2 Background and related work
Neural operators. Learning complex physical systems directly from data is crucial in many scientific and engineering applications (Ghaboussi et al., 1998; Liu et al., 2024c; Ghaboussi et al., 1991; Carleo et al., 2019; Karniadakis et al., 2021; Zhang et al., 2018; Cai et al., 2022; Pfau et al., 2020; He et al., 2021; Jafarzadeh et al., 2024). Often, the governing laws of these systems are unknown and must be inferred from data through learning models that are both resolution-invariant to ensure consistent performance across various levels of discretizations and interpretable to domain experts. Neural operators (NOs) address the first aim by learning mappings between infinite-dimensional function spaces (Li et al., 2020a, b, c; You et al., 2022a; Ong et al., 2022; Cao, 2021; Lu et al., 2019, 2021; Goswami et al., 2022; Gupta et al., 2021). By capturing the relationship between spatial and spatio-temporal data, NOs offer a promising approach to discovering continuum physical surrogates, enabling more accurate and consistent predictions of the underlying responses of complex systems. Nevertheless, NOs lack an interpretable representation of the physical states driving the system, and as a result offer limited insight into the underlying physical mechanisms.
@@
Forward and inverse PDE learning. Existing NOs focus on providing efficient surrogates for the underlying physical systems as forward PDE solvers. They are often employed as black-box universal approximators but lack interpretability of the underlying physical principles. On the other hand, several deep learning methods have been developed as inverse PDE solvers (Fan and Ying, 2023; Molinaro et al., 2023; Jiang et al., 2022; Chen et al., 2023), aiming to reconstruct the PDE parameters from solution data. However, the inverse problem is typically more challenging due to its ill-posedness. To tackle this pathology, many NOs incorporate prior information via governing PDEs (Yang et al., 2021; Li et al., 2021), regularizers (Dittmer et al., 2020; Obmann et al., 2020; Ding et al., 2022; Chen et al., 2023), or additional operator structures (Uhlmann, 2009; Lai et al., 2019; Yilmaz, 2001). However, all these work rely on prior information on the operator form, which is not realistic in many real-world problems where neither the parameter field nor the model form are available. Worth noting is a recent work that leverages the attention mechanism to build a nonlinear kernel map in NO (Yu et al., 2024), with which a data-dependent kernel can be constructed to characterize the inverse mapping from data to the hidden governing mechanism. While effective, this approach does not disentangle the learned kernel or extract physically interpretable parameters. To the best of our knowledge, our proposed DisentangO is the first that tackles both forward (i.e., physics prediction) and inverse (i.e., physics discovery) PDE learning problems, while simultaneously inferring the governing distinct physical parameters from the learned NO.
@@
Disentangled representation learning. Disentangled representation learning seeks to separate data into distinct, interpretable factors, where each factor captures an independent aspect of variation in the input. This area has critical implications for transfer learning, generative modeling, and fairness in AI, as it allows models to generalize across tasks by leveraging isolated features, making them more interpretable and modular. Disentangled representation learning has seen significant advances, primarily driven by models like $\beta$-VAE (Higgins et al., 2017), FactorVAE (Kim and Mnih, 2018), and InfoGAN (Chen et al., 2016), who encourage disentanglement by controlling the trade-off between reconstruction accuracy and latent regularization, or introduce mutual information constraints to encourage the discovery of interpretable latent factors. Early work in disentanglement largely focus on supervised or semi-supervised approaches (Ridgeway and Mozer, 2018; Shu et al., 2019; Mathieu et al., 2019), requiring labeled data for factors of variation. However, more recent advancements have concentrated on unsupervised learning methods (Duan et al., 2019), where disentanglement is achieved without explicit labels. Despite these efforts, fully unsupervised disentanglement remains challenging (Locatello et al., 2019), with ongoing research focusing on improving metrics, robustness, and applicability to more complex and high-dimensional data. In this context, although a large body of research has been done in disentanglement, its applications remain largely confined to fields like computer vision and robotics, where the latent dimensions typically correspond to human-interpretable (often visually recognizable) features, with very limited exploration in areas such as physical system learning. Moreover, existing work focuses on disentangling from data, whereas our work is the first attempt that explores disentanglement from black-box neural network parameters.
3 DisentangO: disentangled representation learning from neural operator parameters
@@
$\mathcal{K}_{\bm{b}}[\bm{u}](\bm{x})=\bm{f}(\bm{x})\text{ ,}\quad\bm{x}\in{\Omega}\text{ .}$ (3.1)
Here, ${\Omega}\subset\mathbb{R}^{s}$ is the domain of interest, $\bm{f}(\bm{x})$ is a function representing the loading on ${\Omega}$, and $\bm{u}(\bm{x})$ is the corresponding solution of this system. $\mathcal{K}_{\bm{b}}$ represents the unknown governing law, e.g., balance laws, which is determined by the (possibly unknown and high-dimensional) parameter field $\bm{b}$. For instance, in a material modeling problem, $\mathcal{K}_{\bm{b}}$ often stands for the constitutive law and $\bm{b}$ can be a vector ($\bm{b}\in\mathbb{R}^{d_{b}}$) representing the homogenized material parameter field or a vector-valued function ($\bm{b}\in L^{\infty}({\Omega};\mathbb{R}^{d_{b}})$) representing the heterogeneous material properties. Both scenarios are considered in our empirical experiments in Section 4.
Many physical modeling tasks can be formulated as either a forward or an inverse PDE-solving problem. In a forward problem setting, one aims to find the PDE solution when given PDE information, including coefficient functions, boundary conditions, initial conditions, and loading sources. That means, given the governing operators $\mathcal{K}$, the parameter (field) $\bm{b}$, and loading field $\bm{f}(\bm{x})$ in equation 3.1, the goal is to solve for the corresponding solution field $\bm{u}(\bm{x})$, through classical PDE solvers (Brenner and Scott, 2007) or data-driven approaches (Lu et al., 2019; Li et al., 2020c; You et al., 2022b). As a result, a forward map is constructed:
@@
Unfortunately, solving an inverse problem is typically more challenging due to the ill-posed nature of the PDE model. In general, a small number of function pairs $(\bm{u},\bm{f})$ from a single system does not suffice in inferring the underlying parameter field $\bm{b}$, making the inverse problem generally non-identifiable (Molinaro et al., 2023). Herein, we propose a novel neural architecture to alleviate the curse of ill-posedness without requiring prior knowledge. The key ingredients are: 1) the construction of a multi-task NO architecture, which solves both the forward (physics prediction) and inverse (physics discovery) problems simultaneously from multiple systems, and 2) a generative model to disentangle the features from NO parameters that contain critical information of $\bm{b}$.
3.1 Variational hierarchical neural operator as a multi-task solver
Figure 1: Illustration of DisentangO’s architecture.
Formally, we consider $S$ training datasets from different systems, each of which contains $n_{train}$ function pairs $(\bm{u},\bm{f})\in\mathcal{U}\times\mathcal{F}$:
$\mathcal{D}_{\rm tr}=\{\{(\bm{u}^{\eta}_{i}(\bm{x}),\bm{f}^{\eta}_{i}(\bm{x}))\}_{i=1}^{n_{train}}\}_{\eta=1}^{S}\text{ .}$ (3.4)
The index $\eta$ corresponds to different hidden parameter field/set $\bm{b}^{\eta}\in\mathcal{B}$. Here, $\mathcal{U}\subset L^{2}({\Omega};\mathbb{R}^{d_{u}})$, $\mathcal{F}\subset L^{2}({\Omega};\mathbb{R}^{d_{f}})$ represent the (infinite-dimensional) Banach spaces of the solution function $\bm{u}$ and loading function $\bm{f}$, respectively, and $\mathcal{B}$ is a finite-dimensional manifold where the hidden parameter $\bm{b}$ takes values on. We further assume that the data satisfies the following generating process:
$\bm{b}|\bm{z}=m(\bm{z})\text{ ,}\;\bm{z}\sim\mathbb{P}_{z}\text{ ,}$ (3.5)
where $\bm{z}\in\mathcal{Z}\subset\mathbb{R}^{d_{z}}$ is the embedding of $\bm{b}$ in the latent space and $m:\mathcal{Z}\rightarrow\mathcal{B}$ is a one-to-one function. Then, we assume for each $\bm{b}^{\eta}$ that there exists an underlying parametric PDE system of equation 3.1. Hence, each function pair $(\bm{u}_{i}^{\eta},\bm{f}^{\eta}_{i})$ satisfies:
$\bm{u}|\bm{f}=\mathcal{K}_{\bm{b}}^{-1}[\bm{f}]\text{ ,}\;\bm{f}\sim\mathbb{P}_{f}\text{ ,}$ (3.6)
where $\mathcal{K}_{\bm{b}}^{-1}$ is the ground-truth solution operator of the parametric PDE system of equation 3.1 with a given parameter $\bm{b}$.
Our first objective is to construct a surrogate operator of $\mathcal{K}_{\bm{b}}^{-1}:\mathcal{F}\rightarrow\mathcal{U}$, as the forward PDE solver for each system:
$\mathcal{G}[\bm{f}^{\eta}_{i};\theta^{\eta}](\bm{x})\approx\mathcal{K}_{\bm{b}^{\eta}}^{-1}[\bm{f}_{i}^{\eta}](\bm{x})=\bm{u}_{i}^{\eta}(\bm{x})\text{ ,}$ (3.7)
where $\theta^{\eta}$ denotes the corresponding trainable parameter set. Then, as the second objective, we construct an inverse map by discovering the hidden parameter field/set from $\theta^{\eta}$:
$\mathcal{H}(\theta^{\eta};\Theta_{H})\approx\tilde{\bm{z}}^{\eta}\text{ .}$ (3.8)
Here, $\Theta_{H}$ is the trainable parameters of $\mathcal{H}$, and $\tilde{\bm{z}}^{\eta}\in\mathbb{R}^{d_{z}}$ denotes the learned latent variables, which can be transformed to the ground-truth latent variable $\bm{z}^{\eta}$ via an invertible function $h$. The goal of disentanglement is not to discover $\bm{b}^{\eta}$, since some features may not be identifiable. For instance, in a Dirichlet boundary condition problem, $\bm{u}^{\eta}_{i}(\bm{x})=0$ for all $\bm{x}\in\partial\Omega$, making $\bm{b}^{\eta}$ not learnable on $\partial\Omega$. Hence, a more realistic goal is to discover the underlying mechanism of $\bm{b}$ in the space of identifiability, i.e., $\bm{z}$, or its equivalent transformation $\tilde{\bm{z}}$.
Note that in the first objective of equation 3.7, one is interested in finding the solution operator $\mathcal{G}^{\eta}$ under a supervised setting, in the form of function-to-function mappings, for all (hidden) parameters $\bm{b}^{\eta}$ in the range of interest; while in the second objective of equation 3.8, the goal is to construct a vector-to-vector mapping from $\theta^{\eta}$ to $\tilde{\bm{z}}^{\eta}$, which is unsupervised due to the hidden physics nature. We propose to employ variational autoencoder (VAE) (Kingma and Welling, 2013) as the representation learning approach for the second objective, and a meta-learned NO architecture as a universal solution operator for the first objective. The key is to pair these architectures as a hierarchical neural operator, which simultaneously solves both forward and inverse modeling problems.
We now define the learning architecture. The overall objective is to maximize the log data likelihood:
$\max E(\log(p(\theta,\mathcal{D})))=\max\left[E(\log(p(\bm{u}|\bm{f},\theta)+\log(p(\theta|\bm{f}))+\log(p(\bm{f}))\right].$

### v2

@@ Abstract
Neural operators (NOs) excel at learning mappings between function spaces, serving as efficient forward solution approximators for PDE-governed systems. However, as black-box solvers, they offer limited insight into the underlying physical mechanism, due to the lack of interpretable representations of the physical parameters that drive the system. To tackle this challenge, we propose a new paradigm for learning disentangled representations from NO parameters, thereby effectively solving an inverse problem. Specifically, we introduce DisentangO, a novel hyper-neural operator architecture designed to unveil and disentangle latent physical factors of variation embedded within the black-box neural operator parameters. At the core of DisentangO is a multi-task NO architecture that distills the varying parameters of the governing PDE through a task-wise adaptive layer, alongside a variational autoencoder that disentangles these variations into identifiable latent factors. By learning these disentangled representations, DisentangO not only enhances physical interpretability but also enables more robust generalization across diverse systems. Empirical evaluations across supervised, semi-supervised, and unsupervised learning contexts show that DisentangO effectively extracts meaningful and interpretable latent features, bridging the gap between predictive performance and physical understanding in neural operator frameworks. Our code and data accompanying this paper are available at https://github.com/ningliu-iga/DisentangO.
@@ 1 Introduction
Interpretability in machine learning (ML) refers to the ability to understand and explain how models make decisions (Rudin et al., 2022; Molnar, 2020). As ML systems grow complex, especially with the use of deep learning and ensemble methods (Sagi & Rokach, 2018), the reasoning behind their predictions can become opaque. Interpretability addresses this challenge by making the models more transparent, enabling users to trust the outcomes, detect biases, and identify potential flaws. It is a critical factor in applying AI responsibly, especially in fields where accountability and fairness are essential (Cooper et al., 2022). In physics, where models endeavor to capture the governing laws and physical principles, understanding how a model arrives at its predictions is vital for verifying that it aligns with known scientific theories. This transparency is key for advancing scientific discovery, validating results, and enhancing trust in models for complex physical systems.
@@
Discovering interpretable representations of physical parameters in learning physical systems is challenging due to the intricate nature of real-world phenomena and the often implicit relationships between variables. In physics, quantities like force, energy, and velocity are governed by well-established laws, and extracting them in a way that aligns with physical intuition requires models that go beyond mere pattern recognition. Traditional ML models may fit the data but fail to provide a physically interpretable way. To address this, recent developments include integrating physical constraints into the learning process (Raissi et al., 2019), such as embedding conservation laws (Liu et al., 2023; 2024a), symmetries (Mattheakis et al., 2019), or invariances (Cohen & Welling, 2016) directly into model architectures. Additionally, methods like symbolic regression (Biggio et al., 2021) and sparse modeling (Carroll et al., 2009) aim to discover simple, interpretable expressions that capture the underlying dynamics. However, balancing model expressivity with interpretability remains a significant challenge, as overly complex models may obscure the true physical relationships, while oversimplified models risk losing critical details of the system’s behavior.
@@
We introduce DisentangO, a novel variational hyper-neural operator architecture to disentangle physical factors of variation from black-box neural operator parameters for solving parametric PDEs. Neural operators (NOs) (Li et al., 2020a; c) learn mappings between infinite-dimensional function spaces in the form of integral operators, making them powerful tools for discovering continuum physical laws by manifesting the mappings between spatial and/or spatiotemporal data; see You et al. (2022a); Liu et al. (2024a; b; 2023); Ong et al. (2022); Cao (2021); Lu et al. (2019; 2021); Goswami et al. (2022); Gupta et al. (2021) and references therein. However, most NOs serve as efficient forward surrogates for the underlying physical system under a supervised learning setting. As a result, they act as black-box universal approximators for a single physical system governed by a fixed set of PDE parameters, and lack interpretability with respect to the underlying physical laws. In contrast, the key innovation of DisentangO lies in the use of a hypernetwork architecture that distills the varying physical parameters of the governing PDE from multiple physical systems through an unsupervised learning setting. The distilled variables are further disentangled into distinct physical factors to enhance physical understanding and promote robust generalization. Consequently, DisentangO effectively extracts meaningful and interpretable physical features, thereby simultaneously solving both the forward and inverse problems. Our key contributions are:
@@
  • We propose a novel variational hyper-neural operator architecture, which we coin DisentangO. DisentangO extracts the key physical factors of variation from black-box neural operator parameters of multiple physical systems. These factors are then disentangled into distinct latent factors that enhance physical interpretation and generalization across different physical systems.
@@
  • We explore the practical utility of disentanglement and perform experiments across a broad range of settings including supervised, semi-supervised, and unsupervised learning. Results show that DisentangO effectively extracts meaningful physical features.
@@ 2 Background and related work
Neural operators. Learning complex physical systems from data is essential in many scientific applications (Carleo et al., 2019; Liu et al., 2024c; Karniadakis et al., 2021; Zhang et al., 2018; Cai et al., 2022; Pfau et al., 2020; He et al., 2021; Jafarzadeh et al., 2024). When governing laws are unknown, models must be both resolution-invariant for consistent performance across discretizations and interpretable for domain experts. Neural operators (NOs) achieve the former by learning mappings between infinite-dimensional function spaces (Li et al., 2020a; b; c; You et al., 2022a; Ong et al., 2022; Cao, 2021; Lu et al., 2019; 2021; Goswami et al., 2022; Gupta et al., 2021), enabling accurate and consistent predictions of continuum physical surrogates. However, NOs lack interpretable representations of physical states, limiting their ability to reveal underlying physical mechanisms.
Hypernetworks. Hypernetworks (Ha et al., 2016; Chauhan et al., 2024) are a class of neural network (NN) architectures that use one NN to generate weights for another NN, both trained in an end-to-end manner. This allows soft weight sharing across tasks, benefiting transfer learning and dynamic information sharing (Chauhan et al., 2023). Hypernetworks can also be employed as a versatile technique in existing NN architectures. For instance, in Nguyen et al. ; Oh & Peng (2022), a hypernetwork is employed to generate parameters for a VAE model and enable multi-task learning. Similarly, Lee et al. (2023) integrates the hypernetwork architecture with neural operators. However, none of the existing work discusses the capability of hypernetworks in hidden physics discovery.
@@
Forward and inverse PDE learning. Existing NOs serve as efficient surrogates for forward PDE solving but often act as black-box approximators, lacking interpretability. In contrast, deep learning methods for inverse PDE solving (Fan & Ying, 2023; Molinaro et al., 2023; Jiang et al., 2022; Chen et al., 2023) aim to reconstruct PDE parameters from solution data but face challenges due to ill-posedness. To address this, many NOs incorporate prior information via governing PDEs (Yang et al., 2021; Li et al., 2021), regularizers (Dittmer et al., 2020; Obmann et al., 2020; Ding et al., 2022; Chen et al., 2023), or structured operators (Lai et al., 2019; Yilmaz, 2001). However, these methods assume prior knowledge of the model form, which is often unrealistic. A recent approach (Yu et al., 2024) employs attention to construct a data-dependent kernel for inverse mapping but does not disentangle the learned kernel or extract interpretable parameters. To our knowledge, DisentangO is the first to tackle both forward (physics prediction) and inverse (physics discovery) PDE learning while simultaneously identifying distinct physical parameters from the learned NO.
@@
Disentangled representation learning. Disentangled representation learning separates data into distinct, interpretable factors, each capturing an independent variation. It has critical implications for transfer learning, generative modeling, and AI fairness, enabling models to generalize by leveraging isolated features. Key advances have been driven by models like $\beta$-VAE (Higgins et al., 2017), FactorVAE (Kim & Mnih, 2018), and InfoGAN (Chen et al., 2016), which promote disentanglement through latent regularization and mutual information constraints. While early work focused on supervised or semi-supervised methods (Ridgeway & Mozer, 2018; Shu et al., 2019; Mathieu et al., 2019), recent efforts target unsupervised learning (Duan et al., 2019), though challenges remain (Locatello et al., 2019), with ongoing efforts to improve metrics, robustness, and applicability to complex data. Most research has centered on computer vision and robotics, where latent factors have human-interpretable visual meanings, while its exploration in physical system learning remains limited (Lingsch et al., 2024; Tong et al., 2024; Fotiadis et al., 2023). Moreover, existing work disentangles from data, whereas our work is the first to disentangle from black-box NN parameters.
3 DisentangO
Figure 1: Overview of the DisentangO architecture. Each task correspond to a different (hidden) PDE parameter $\bm{b}$. For illustration, the same input function $f^{\eta}_{i}$ is shown for multiple tasks to highlight that different parameter fields $b^{\eta}$ can produce different output functions $u_{i}^{\eta}$ under identical input $f$; in practice, $f$ may vary across tasks. The task-specific lifting parameters $\theta_{P}^{\eta}$ are encoded and reconstructed through a VAE, and the reconstructed parameters $\hat{\theta}_{P}^{\eta}$ are fed into the iterative Fourier layers to form the task-specific neural operator $G^{\eta}$. Loss components are overlaid to indicate where each term in the objective $L_{loss}$ is computed.
@@ 3 DisentangO
[…]
Here, ${\Omega}\subset\mathbb{R}^{s}$ is the domain of interest, $\bm{f}(\bm{x})$ is a function representing the loading on $\Omega$}, $\bm{u}(\bm{x})$ is the corresponding solution of this system. $\mathcal{K}_{\bm{b}}$ represents the unknown governing law, e.g., balance laws, determined by the (possibly unknown and high-dimensional) parameter field $\bm{b}$. For instance, in a material modeling problem, $\mathcal{K}_{\bm{b}}$ often stands for the constitutive law and $\bm{b}$ can be a vector ($\bm{b}\in\mathbb{R}^{d_{b}}$) representing the homogenized material parameter field or a vector-valued function ($\bm{b}\in L^{\infty}({\Omega};\mathbb{R}^{d_{b}})$) representing the heterogeneous material properties. Both scenarios are considered in our empirical experiments in Section 4.
Many physical modeling tasks can be formulated as either a forward or an inverse PDE-solving problem. In a forward problem setting, one aims to find the PDE solution when given PDE information, including coefficient functions, boundary conditions, initial conditions, and loading sources. That means, given the governing operators $\mathcal{K}$, the parameter (field) $\bm{b}$, and loading field $\bm{f}(\bm{x})$ in equation 3.1, the goal is to solve for the corresponding solution field $\bm{u}(\bm{x})$, through classical PDE solvers (Brenner & Scott, 2007) or data-driven approaches (Lu et al., 2019; Li et al., 2020c). As a result, a forward map is constructed:
@@
Unfortunately, solving an inverse problem is typically more challenging due to the ill-posed nature of the PDE model. In general, a small number of function pairs $(\bm{u},\bm{f})$ from a single system does not suffice in inferring the underlying parameter field $\bm{b}$, making the inverse problem generally non-identifiable (Molinaro et al., 2023).
Herein, we propose a novel neural architecture to alleviate the curse of ill-posedness without access to the exact governing partial differential equation (PDE). The key ingredients are: 1) the construction of a multi-task NO architecture, which solves both the forward (physics prediction) problems simultaneously from multiple PDE systems with different hidden parameters; 2) a generative model to disentangle the key features from NO parameters that contain critical information of $\bm{b}$, as the inverse (physics discovery) problem solver.
3.1 Variational hyper-neural operator as a multi-task solver
Notation and data model assumptions. We denote by $\bm{f}(\bm{x})$ and $\bm{u}(\bm{x})$ the input and output functions of the NO, respectively. Let $p(\cdot)$ denote a probability density function, $\mathbb{E}(\cdot)$ the expectation, and $\left|\left|\cdot\right|\right|$} the $l^{2}$-norm. We assume (noisy) observations of both $\bm{u}$ and $\bm{f}$ are available on a common physical domain $\Omega$, where $\bm{u}\in\mathcal{U}\subset L^{2}(\Omega;\mathbb{R}^{d_{u}})$ and $\bm{f}\in\mathcal{F}\subset L^{2}(\Omega;\mathbb{R}^{d_{f}})$. Here, $\mathcal{U}$ and $\mathcal{F}$ are the Banach spaces of solution fields and loading fields, respectively. Formally, we consider $S$ training datasets $\mathcal{D}^{\eta}$, $\eta=1,\dots,S$, each corresponding to the same PDE equation 3.1 but with a different (hidden) physical parameter $\bm{b}^{\eta}\in\mathcal{B}$, with $\mathcal{B}$ the parameter space. In our multi-task learning model, each dataset corresponds to one task. We assume the hidden parameter is generated according to:
$\bm{b}\sim\mathbb{P}_{b}\text{ ,}\;\bm{z}\sim p(\bm{z}|\bm{b})\text{ ,}$ (3.4)
where $\bm{z}\in\mathcal{Z}\subset\mathbb{R}^{d_{z}}$ is the latent embedding of $\bm{b}$. Each dataset contains measurements from $n^{\eta}_{train}$ function pairs $\{(\bm{u}_{i}^{\eta}(\bm{x}),\bm{f}_{i}^{\eta}(\bm{x}))\}_{i=1}^{n^{\eta}_{train}}$. Although our method accommodates datasets with varying numbers of function pairs, we use the same number of function pairs in this work for simplicity and denote $n^{\eta}_{train}=n_{train}$. While actual observations are gathered on a discrete sensor set $\chi=\{\bm{x}_{j}\}_{j=1}^{\#\chi}$ and inevitably contain observational noise of the solution, we adopt the standard assumption that this noise is additive and i.i.d. Letting $\mathcal{G}^{\dagger}[\bm{f};\bm{b}]$ denote the true forward operator of equation 3.1, we write the data model associated with multi-task learning as: